OpenStudy (anonymous):

Solve sqrt7x+36 =x+4 x=

OpenStudy (anonymous):

x = 1/2 (71+3 sqrt(105))

OpenStudy (anonymous):

sqrt{7x} = x + 40 7x = (x+40)^2 Quadratic Formula Now, after expanding